# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v6

@@ 5.2 Token-Based Generation
Early Transformer applications.
Even before the advent of LLMs (Brown et al., 2020; Achiam et al., 2023), Transformers have already made a significant impact on time-series modeling. Video Transformer (Weissenborn et al., 2020) pioneer the use of Transformer architectures in video synthesis by introducing an auto-regressive modeling approach. Despite its success, it inherits common limitations of Transformer-based models, including high training resource demands and slow inference speed.
@@
The Latent Video Transformer (LVT) (Rakhimov et al., 2021) introduces a novel latent-space approach that auto-regressively models temporal dynamics and predicts future features, significantly reducing computational overhead. Other works, such as VideoGen (Zhang et al., 2020c) and Video VQ-VAE (Walker et al., 2021), also leverage the VQ-VAE framework (Van Den Oord et al., 2017) to extract discrete tokens for video prediction. In contrast, Phenaki (Villegas et al., 2022) improves upon the ViViT (Arnab et al., 2021) architecture to extract continuous tokens. The NUWA framework (Wu et al., 2022a) proposes a versatile 3D Transformer-based encoder-decoder architecture that is adaptable to diverse data modalities and tasks, further demonstrating the potential of Transformers in video synthesis. NUWA-Infinity (Liang et al., 2022) builds upon this with an innovative generation mechanism designed to enable infinite high-resolution video synthesis, reflecting ongoing efforts to unify generative tasks across modalities.
@@ 6 Application Realms
The applications of FFS span a wide range of domains, underscoring its growing significance across fields.
@@ 7 Related Work
The field of video prediction, along with related areas such as action recognition and spatio-temporal predictive learning, has witnessed significant progress in recent years, largely driven by deep learning techniques. Several comprehensive surveys have provided overviews of state-of-the-art methods, benchmark datasets, and evaluation protocols in this domain. Zhou et al. (2020) review next-frame prediction models developed prior to 2020, categorizing them into sequence-to-one and sequence-to-sequence architectures. The survey compares these approaches by analyzing their architectural designs and loss functions, and provides quantitative performance comparisons based on standard datasets and evaluation metrics. Oprea et al. (2020) present a comprehensive review of deep learning methods for video prediction, outlining fundamental concepts and analyzing existing models based on a proposed taxonomy. The survey also includes experimental results to enable quantitative assessment of the state of the art. Rasouli (2020) provide an overview of vision-based prediction algorithms with a focus on deep learning approaches. They categorize prediction tasks into video prediction, action prediction, trajectory prediction, body motion prediction, and other related applications, and discuss common network architectures, training strategies, data modalities, evaluation metrics, and benchmark datasets. Kong & Fu (2022) survey state-of-the-art techniques in action recognition and prediction, covering existing models, representative algorithms, technical challenges, action datasets, evaluation protocols, and future research directions. Tan et al. (2023) introduce OpenSTL, a unified benchmark for spatio-temporal predictive learning, which categorizes methods into recurrent-based and recurrent-free models. The paper provides standardized evaluations on multiple datasets and offers an in-depth analysis of how model architecture and dataset characteristics influence performance.
@@
Xing et al. (2023) present a comprehensive review of video diffusion models in the era of AI-generated content, categorizing existing works into video generation, video editing, and other video understanding tasks. The survey provides an in-depth analysis of the literature in these areas and discusses current challenges and future research trends. Li et al. (2024) survey recent advances in long video generation, summarizing existing methods into two key paradigms: divide-and-conquer and temporal auto-regressive modeling. They also provide a comprehensive overview and categorization of datasets and evaluation metrics, and discuss emerging challenges and future directions in this rapidly evolving field. Sun et al. (2024b) review Sora, OpenAI’s text-to-video model, categorizing the related literature into three themes—evolutionary generators, excellence in pursuit, and realistic panoramas—while also discussing datasets, evaluation metrics, existing challenges, and future directions. Complementarily,  Liu et al. (2024b) provide a comprehensive analysis of Sora’s underlying technologies, system design, current limitations, and its potential role in the broader landscape of large vision models.  Melnik et al. (2024) offer an in-depth exploration of the critical components of video diffusion models, focusing on their applications, architectural design, and temporal dynamics modeling.
@@ 8 Conclusion
Future prediction models that do not incorporate generative content do not experience significant negative impacts on their performance improvements. With the rise of AI-generation methods, the way we create and consume video content is rapidly evolving. While AI-generated technologies hold the potential to save substantial time for content marketers and video creators, they also raise numerous ethical concerns. On the positive side, AI video generation can save both time and money for content creators, expand their imagination, and democratize video creation, making it easier for small businesses and individuals to compete with companies that have large marketing budgets. However, it’s inevitable that distinguishing between real and AI-generated content will become increasingly challenging. Without proper safeguards, there is a risk of misinformation, propaganda, and the manipulation of public opinion. Generative videos might misrepresent people and events, reflect biases and discrimination from the training set, pose legal and copyright issues, and potentially lead to job losses for human video editors and animators. Apart from imposing constraints on content generation service providers, such as preventing inappropriate generation and adding watermarks, the exploration of AI-Generated Media Detection methods is also very useful. To mitigate risks, content generation service providers can prevent inappropriate generation and add watermarks. Exploring methodologies for detecting AI-generated media (Zou et al., 2025; Chen et al., 2024a) holds significant value as well.

### v7

@@ 5.2 Token-Based Generation
Even before the advent of LLMs (Brown et al., 2020; Achiam et al., 2023), Transformers have already made a significant impact on time-series modeling. Video Transformer (Weissenborn et al., 2020) pioneer the use of Transformer architectures in video synthesis by introducing an auto-regressive modeling approach. Despite its success, it inherits common limitations of Transformer-based models, including high training resource demands and slow inference speed. The Latent Video Transformer (LVT) (Rakhimov et al., 2021) introduces a novel latent-space approach that auto-regressively models temporal dynamics and predicts future features, significantly reducing computational overhead. Other works, such as VideoGen (Zhang et al., 2020c) and Video VQ-VAE (Walker et al., 2021), also leverage the VQ-VAE framework (Van Den Oord et al., 2017) to extract discrete tokens for video prediction. In contrast, Phenaki (Villegas et al., 2022) improves upon the ViViT (Arnab et al., 2021) architecture to extract continuous tokens. The NUWA framework (Wu et al., 2022a) proposes a versatile 3D Transformer-based encoder-decoder architecture that is adaptable to diverse data modalities and tasks, further demonstrating the potential of Transformers in video synthesis. NUWA-Infinity (Liang et al., 2022) builds upon this with an innovative generation mechanism designed to enable infinite high-resolution video synthesis, reflecting ongoing efforts to unify generative tasks across modalities.
@@ 6 Application Realms
The applications of FFS are extensive and diverse, permeating numerous domains and highlighting its increasing importance as a versatile tool. Across these domains, FFS serves as a fundamental enabler for predictive modeling and decision-making, allowing systems to anticipate future states based on past observations. This predictive capability is crucial for enhancing safety, efficiency, and adaptability in dynamic environments.
@@ 7 Related Work
In the related work section, we delineate our survey’s unique focus on FFS by comparing it with existing literature on video prediction and video diffusion models, highlighting both similarities and key distinctions.
@@
The field of video prediction, along with related areas such as action recognition and spatio-temporal predictive learning, has witnessed significant progress in recent years, largely driven by deep learning techniques. Several comprehensive surveys have provided overviews of state-of-the-art methods, benchmark datasets, and evaluation protocols in this domain. Zhou et al. (2020) review next-frame prediction models developed prior to 2020, categorizing them into sequence-to-one and sequence-to-sequence architectures. The survey compares these approaches by analyzing their architectural designs and loss functions, and provides quantitative performance comparisons based on standard datasets and evaluation metrics. Oprea et al. (2020) present a comprehensive review of deep learning methods for video prediction, outlining fundamental concepts and analyzing existing models based on a proposed taxonomy. The survey also includes experimental results to enable quantitative assessment of the state-of-the-art. Rasouli (2020) provide an overview of vision-based prediction algorithms with a focus on deep learning approaches. They categorize prediction tasks into video prediction, action prediction, trajectory prediction, body motion prediction, and other related applications, and discuss common network architectures, training strategies, data modalities, evaluation metrics, and benchmark datasets. While these surveys focus on early technical developments in video prediction, our survey specifically emphasizes the synthesis aspect of FFS and its evolution towards generative methodologies, highlighting the rising significance of generative models.
Kong & Fu (2022) survey state-of-the-art techniques in action recognition and prediction, covering existing models, representative algorithms, technical challenges, action datasets, evaluation protocols, and future research directions. Tan et al. (2023) introduce OpenSTL, a unified benchmark for spatio-temporal predictive learning, which categorizes methods into recurrent-based and recurrent-free models. The paper provides standardized evaluations on multiple datasets and offers an in-depth analysis of how model architecture and dataset characteristics influence performance. For our survey, these reviews are excellent supplements regarding technical details such as architectures and benchmarks.
@@
More recently, the emergence of video diffusion models has led to several specialized surveys.  Xing et al. (2023) present a comprehensive review of video diffusion models in the era of AI-generated content, categorizing existing works into video generation, video editing, and other video understanding tasks. The survey provides an in-depth analysis of the literature in these areas and discusses current challenges and future research trends. Li et al. (2024) survey recent advances in long video generation, summarizing existing methods into two key paradigms: divide-and-conquer and temporal auto-regressive modeling. They also provide a comprehensive overview and categorization of datasets and evaluation metrics, and discuss emerging challenges and future directions in this rapidly evolving field. Sun et al. (2024b) review Sora, OpenAI’s text-to-video model, categorizing the related literature into three themes—evolutionary generators, excellence in pursuit, and realistic panoramas—while also discussing datasets, evaluation metrics, existing challenges, and future directions. Complementarily,  Liu et al. (2024b) provide a comprehensive analysis of Sora’s underlying technologies, system design, current limitations, and its potential role in the broader landscape of large vision models.  Melnik et al. (2024) offer an in-depth exploration of the critical components of video diffusion models, focusing on their applications, architectural design, and temporal dynamics modeling. These surveys offer highly specific analyses of recent models. In contrast, our survey provides a broader review of the FFS methodology, positioning models like Sora as examples within the evolving landscape of the generative approaches to FFS.
@@ 8 Conclusion
The performance improvements of future prediction models that do not incorporate generative content do not bring significant negative social impacts. With the rise of AI-generation methods, the way we create and consume video content is rapidly evolving. While AI-generated technologies hold the potential to save substantial time for content marketers and video creators, they also raise numerous ethical concerns. On the positive side, AI video generation can save both time and money for content creators, expand their imagination, and democratize video creation, making it easier for small businesses and individuals to compete with companies that have large marketing budgets. However, it’s inevitable that distinguishing between real and AI-generated content will become increasingly challenging. Without proper safeguards, there is a risk of misinformation, propaganda, and the manipulation of public opinion. Generative videos might misrepresent people and events, reflect biases and discrimination from the training set, pose legal and copyright issues, and potentially lead to job losses for human video editors and animators. Apart from imposing constraints on content generation service providers, such as preventing inappropriate generation and adding watermarks, the exploration of AI-Generated Media Detection methods is also very useful. To mitigate risks, content generation service providers can prevent inappropriate generation and add watermarks. Exploring methodologies for detecting AI-generated media (Zou et al., 2025; Chen et al., 2024b) holds significant value as well.